OpenStudy (anonymous):

I really need help with this Linear Equation. The number of gallons of bottled water N consumed by a person in on year is approximated by N=0.733t+8.398, where t is the number of years since 1971. Predict the number of gallons consumed by a person in 1978.

OpenStudy (anonymous):

do you know the equation of line?

OpenStudy (anonymous):

if you notice N = 0.733t+8.398 is the same as y = mx+b

OpenStudy (anonymous):

Yes I know i'm on mymathlab and it's so confusing to me how they changed it abit for this problem. I'm not looking for a direct answer just need it explained to me. I'm sorry. :(

OpenStudy (anonymous):

it's okay. as you study, you get familiarize with the equation. so, it is said that 't' is from 1971, and predict how many gallons consumed in 1978. you just have to get the interval, t=1978-1971. so t=7.. substitute 7 to the equation. we can get the the number of gallons consumed for 7 years.

OpenStudy (anonymous):

wow i'm so dumb. they didn't even explain that too me in the example thank you so much! this really helped me alot! much love

OpenStudy (anonymous):

it is intended not to, for you to practice. :D anyways, you're welcome! :)